If a rectangle has an area of 100 square inches and a perimeter of 50 inches, what are its dimensions?
Answer by TimothyLamb(4379)

Lw = 100
L = 100/w
2L + 2w = 50
---
2L + 2w = 50
2(100/w) + 2w = 50
200/w + 2w = 50
200w/w + 2ww = 50w
2w^2 - 50w + 200 = 0
---
the above quadratic equation is in standard form, with a=2, b=-50, and c=200
---
to solve the quadratic equation, plug this:
2 -50 200
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
w = 5
w = 20
---
both roots will solve the problem, but traditionally width < length, so choose w = 5.
---
Answer:
w = 5 in
L = 20 in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php